1) Eight times the sum of a number and -2 is the same as nine times the numbers. Find the number
2) Twice the difference of a number and seven is five times the number plus one. Find the number
3) If the sum of a number and -3 is doubled the results is 1/2 times the number. Find the number

Answer by ewatrrr(24785)


Hi,
*Note; Key here is to write as we read after naming the variable
1) Let x represent the number
8(x-2) = 9x
2)Let x represent the number
2(x-7) = 5x + 1
3)Let x represent the number
2[x +(-3)] = x/2
solve the above for x and check your answer